# The Giroux Correspondence in dimension 3

### v1

@@ 3.2. Bridge Heegaard splittings
Figure 6. Left: Adding successive bypasses produces a sequence of (red) handlbodies $U_{i}$ that are homeomorphic but not weakly contact isotopic to $\nu(K_{U})$. Right: Adding the $1$-handles produces a bridge splitting $\widehat{H}=(\widehat{\Sigma},\widehat{U},\widehat{V})$; here $\widehat{U}$ is indicated in blue.
Begin with a smooth Heegaard splitting $\mathcal{H}=(\Sigma,U,V)$. Since the complement $X=M\setminus\big{(}\nu(K_{U})\cup\nu(K_{V})\big{)}$ is topologically a product, Theorem 2.6 implies that up to weak contact isotopy, $X$ decomposes as a sequence of bypasses $X_{1}\ast\dots\ast X_{f}$ attached to $\partial\nu(K_{U})$.111the triple $(\nu(K_{U}),\nu(K_{V})),X)$ is called a ”patty” in [HBH]. Attaching only the first $i$ bypasses in the sequence determines a new handlebody $U_{i}$ with convex boundary; letting $i$ range over the indices of all bypasses produces a sequence of smoothly isotopic Heegaard surfaces $\Sigma_{i}:=\partial U_{i}$. Each $\Sigma_{i}$ has a bypass arc $c_{i+1}$ that determines the bypass $X_{i+1}$. Isotopy of $c_{i}$ through bypass arcs preserves the weak contact isotopy class of $U_{i}$, so we may push each bypass arc off the faces of any $2$-handles already attached.
@@
^$\widehat{\mathcal{H}}(X_{1}\ast\dots\ast X_{f})=\big{(}\partial\widehat{U},\widehat{U},M\setminus\text{int}(\widehat{U})\big{)}.$
Attaching a contact $1$-handle preserves tightness, so it is immediate that $\widehat{U}$ is tight. The complementary $\widehat{V}$ is also tight, as turning the manifold upside down builds $\widehat{V}$ from $1$-handles attached to $\nu(K_{V})$. In fact, the splitting ^$\widehat{\mathcal{H}}(X_{1}\ast\dots\ast X_{f})=(\widehat{\Sigma},\widehat{U},\widehat{V})$ is convex, as co-core discs for each $1$-handle, together with meridional discs for the neighbourhoods of the skeleta, form a complete set of product discs.
@@
Although the term “bridge” is new, this construction appears in both [HBH] and [LV] as a technique to construct a convex Heegaard splitting from an arbitrary smooth Heegaard splitting.
When the original Heegaard surface $\Sigma$ is already convex in the splitting $\mathcal{H}=(\Sigma,U,V)$, one may consider the relative bridge ^$\widehat{\mathcal{H}}^{U}$ built from a bypass decomposition of $M\setminus\big{(}\nu(K_{U})\cup V\big{)}$, although this need not be convex. Similarly, ^$\widehat{\mathcal{H}}^{V}$ is constructed from a bypass decomposition of $M\setminus\big{(}U\cup\nu(K_{V})\big{)}$. The relative bridges will be used in Proposition 3.10.
@@
^$\widehat{\mathcal{H}}(X_{1}\ast\dots\ast X_{f})=(\widehat{\Sigma},\widehat{U},\widehat{V})$
@@
In order to construct a bridge splitting from a bypass decomposition, all the bypass attaching arcs must be isotoped off the faces of an $2$-handles already attached, but there is no canonical way to perform this isotopy. Although different isotopies preserve the weak contact isotopy type of each $U_{i}$, they may produce handlebodies which are inequivalent when the 2-handles are removed. We will show that an isotopy of the attaching 00-sphere for a $1$-handle on $\Sigma_{i}$ preserves the positive stabilisation class of the bridge $\widehat{\mathcal{H}}$.
Figure 7. Left: Before constructing the bridge of $\mathcal{H}$, bypass arcs must be isotoped off the faces of the $2$-handles. Centre: Isotoping $c$ across the face of a $2$-handle on $\Sigma_{i}$ preserves the weak contact isotopy class of $U_{i}$. Right: When $h_{i}^{2}$ is removed, the bypass arcs $c^{R}$ and $c^{L}$ are not isotopic on $\widehat{\Sigma}$.
@@
Figure 9 shows that attaching the bypass along $c_{T}$ onto $\widehat{U}^{L}$ produces a handlebody weakly contact isotopic to $\widehat{U}^{R}$. Then by Theorems 2.5 and 3.4, the convex Heegaard decompositions ^$\widehat{\mathcal{H}}^{L}$ and ^$\widehat{\mathcal{H}}^{R}$ admit a common positve stabilisation. ∎
Proposition 3.8 implies that up to positive stabilisation, the bridge splitting ^$\widehat{\mathcal{H}}=(\widehat{\Sigma},\widehat{U},\widehat{V})$ depends only on $K_{U}$, $K_{V}$, and the bypass decomposition $X=X_{1}\ast\cdots\ast X_{f}$. As we show next, the bridge $\widehat{\mathcal{H}}$ is actually independent – again, up to positive stabilisation – of the bypass decomposition.
@@
Given a smooth Heegaard splitting $\mathcal{H}=(\Sigma,U,V)$ let $K_{U}$ and $K_{V}$ be Legendrian skeletons for $U$ and $V$ and let $X=M\setminus\big{(}\nu(K_{U})\cup\nu(K_{V})\big{)}$. Let $X_{1}\ast\dots\ast X_{f}$ and $X^{\prime}_{1}\ast\dots\ast X^{\prime}_{f^{\prime}}$ be distinct bypass decompositions of $X$, and let ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(X_{1}\ast\dots\ast X_{f})$ and ^=^$\widehat{\mathcal{H}}^{\prime}=\widehat{\mathcal{H}}^{\prime}(X^{\prime}_{1}\ast\dots\ast X^{\prime}_{f})$ be the associated bridge Heegaard splittings formed by attaching only the $1$-handles of the respective bypasses to $\nu(K_{U})$. Then $\widehat{\mathcal{H}}$ and $\widehat{\mathcal{H}}^{\prime}$ admit a common positive stabilisation.
@@
Fix a bypass decomposition of $X$. As above, let ^$\widehat{\mathcal{H}}=(\widehat{\Sigma},\widehat{U},\widehat{V})$ be the Heegaard splitting formed by attaching the associated $1$-handles to $\nu(K_{U})$. Now consider an alternative bypass decomposition that differs by the insertion of a trivial bypass $X_{T}=h^{1}_{T}\cup h^{2}_{T}$ attached along a trivial bypass arc $c_{T}$ somewhere in the sequence, and let ^$\widehat{\mathcal{H}}_{T}=(\widehat{\Sigma}_{T},\widehat{U}_{T},\widehat{V}_{T})$ denote the associated Heegaard splitting. Observe that the genus of ^$\widehat{\mathcal{H}}_{T}$ is larger than the genus of $\widehat{\mathcal{H}}$ because the additional $1$-handle $h_{T}^{1}$ is attached.
We claim that regardless of where the trivial bypass was inserted, $\widehat{\mathcal{H}}$ and ^$\widehat{\mathcal{H}}_{T}$ admit a common positive stabilisation. Recall that Proposition 3.8 allows us to isotope each attaching sphere along respective components of $\Gamma_{\Sigma_{i}}$ while preserving the positive stabilisation class of the bridge splitting. Using this flexibility, ensure that no handle attached after $h_{T}^{1}$ has an attaching sphere that intersects the disc supporting the trivial bypass. Not only can any attaching sphere be isotoped off this region, but we may also impose the stronger condition that subsequent attaching spheres are disjoint from the larger disc shown in Figure 10 that contains the bigon certifying $c_{T}$ as a trivial bypass arc. As a consequence, $c_{T}$ remains a trivial bypass arc on $\widehat{\Sigma}$; we will use this below. It follows that we may apply a sequence of Far Commutation moves to get a new bypass decomposition in which the bypass along $c_{T}$ is attached last.
@@
Given a tight Heegaard splitting $\mathcal{H}=(\Sigma,U,V)$, let $K_{U}$, $K_{V}$ be Legendrian skeletons for $U$ and $V$, respectively. Let $X=X_{1}\ast\cdots\ast X_{f}$ be a bypass decomposition of the complement of $\nu(K_{V})$ in $V$, where the index is the order in which the bypass slices are attached to $U$. Let ^$\widehat{\mathcal{H}}^{U}$ denote the relative bridge splitting with handlebodies $\widehat{U}=U\cup\big{(}\bigcup_{i=1}^{f}h^{1}_{i}\big{)}$ and $\widehat{V}=M\setminus\text{int}(\widehat{U})$. Then the refinements of the Heegaard splittings ^$\widehat{\mathcal{H}}^{U}$ and $\mathcal{H}$ admit a common positive stabilisation.
The relative bridge splitting ^$\widehat{\mathcal{H}}^{U}$ is tight by construction, so one may flip this Heegaard splitting upside down and apply Proposition 3.10 again to get the following consequence:
@@
The proof of Proposition 3.10 is essentially the same argument as the proof of [LV, Theorem 6.1.]. Since it was not stated in full generality there, we indicate how it follows from the techniques of [LV].
@@
Define the sequence of intermediate Heegaard splittings ^$\widehat{\mathcal{H}}_{j}=(\widehat{\Sigma}_{j},\widehat{U}_{j},\widehat{V}_{j})$ built by attaching only the first $j$ of the $1$-handles:
@@
Beginning with the case $j=1$, construct the refinement ~$\widetilde{\mathcal{H}}=(\widetilde{\Sigma},\widetilde{U},\widetilde{V})$ of $\mathcal{H}$.
Recall that the refinement is constructed by drilling tunnels through meridional discs for the handlebodies. We may choose these discs to be disjoint from the bypass $X_{1}=h^{1}_{1}\cup h^{2}_{1}$, so adding the $1$-handle $h_{1}^{1}$ to $U$ commutes with this drilling. Thus attaching $h^{1}_{1}$ to the refined $\widetilde{U}$ produces the same Heegaard splitting as taking the refinement of ^$\widehat{\mathcal{H}}_{1}$. As noted in Section 3.1, attaching $h_{1}^{1}$ in to $\widetilde{U}$ in the convex refinement is a positive stabilisation, which completes the base case.
Now proceed by induction on $j$. For the inductive step, note that if ^$\widehat{\mathcal{H}}_{j}$ is tight, then ^$\widehat{\mathcal{H}}_{j+1}$ is also tight because $\widehat{V}_{j+1}$ is a proper subset of the tight $\widehat{V}_{j}$, while $\widehat{U}_{j+1}$ is built from the tight $\widehat{U}_{j}$ by adding a contact $1$-handle. The argument above shows that refining and then adding a $1$-handle produces the same splitting as first adding the $1$-handle and then refining, which preserves the positive stabilisation class. This establishes the inductive step. ∎
@@ 3.3. Proof of the Giroux Correspondence
First, we repeat an argument from the proof of Theorem 6.1 in [LV] that lets us restrict to Heegaard splittings that are smoothly isotopic:
@@
Choose a Legendrian bouquet $K_{*}$ as a skeleton for each handlebody, where $*\in\{U,U^{\prime},V,V^{\prime}\}$. Since $K_{{U^{\prime}}}\cup K_{{V^{\prime}}}$ is topologically isotopic to $K_{{U}}\cup K_{{V}}$, some stabilisations of these bouquets are Legendrian isotopic. As we may stabilise each $K_{*}$ while preserving it as a skeleton of the respective handlebody, suppose now that the skeletons $K_{{U^{\prime}}}\cup K_{{V^{\prime}}}$ are Legendrian isotopic to $K_{{U}}\cup K_{{V}}$.
@@
Concatenating these produces a bypass decomposition of $X=M\setminus\big{(}\nu(K_{U})\cup\nu(K_{V})\big{)}$. As usual, let $\widehat{U}$ denote the handlebody formed by attaching all of the $1$-handles to $\nu(K_{U})$. Setting ^$\widehat{\mathcal{H}}=(\partial\widehat{U},\widehat{U},M\setminus\text{int}\widehat{U})$, Corollary 3.11 implies that the refinement $\widetilde{\mathcal{H}}$ and the bridge $\widehat{\mathcal{H}}$ admit a common positive stabilisation. To complete the argument, we recall Remark 3.5: since $\mathcal{H}$ is convex, any refinement of $\widetilde{\mathcal{H}}$ is a positive stabilisation of $\mathcal{H}$ itself. Thus $\mathcal{H}$ and $\widehat{\mathcal{H}}$ admit a common positive stabilisation.

### v2

@@ 3.2. Bridge Heegaard splittings
[…]
Begin with a smooth Heegaard splitting $\mathcal{H}=(\Sigma,U,V)$. Since the complement $X=M\setminus\big(\nu(K_{U})\cup\nu(K_{V})\big)$ is topologically a product, Theorem 2.8 implies that up to weak contact isotopy, $X$ decomposes as a sequence of bypasses $X_{1}\ast\dots\ast X_{f}$ attached to $\partial\nu(K_{U})$.111the triple $(\nu(K_{U}),\nu(K_{V})),X)$ is called a ”patty” in [10]. Attaching only the first $i$ bypasses in the sequence determines a new handlebody $U_{i}$ with convex boundary; letting $i$ range over the indices of all bypasses produces a sequence of smoothly isotopic Heegaard surfaces $\Sigma_{i}:=\partial U_{i}$. Each $\Sigma_{i}$ has a bypass arc $c_{i+1}$ that determines the bypass $X_{i+1}$. Isotopy of $c_{i}$ through bypass arcs preserves the weak contact isotopy class of $U_{i}$, so we may push each bypass arc off the faces of any $2$-handles already attached.
@@
$).\widehat{\mathcal{H}}(X_{1}\ast\dots\ast X_{f})=\big(\partial\widehat{U},\widehat{U},M\setminus\text{int}(\widehat{U}$$\big)$.
[…]
@@
Although the term “bridge” is new, this construction appears in both [10] and [16] as a technique to construct a convex Heegaard splitting from an arbitrary smooth Heegaard splitting.
When the original Heegaard surface $\Sigma$ is already convex in the splitting $\mathcal{H}=(\Sigma,U,V)$, one may consider the relative bridge $\widehat{\mathcal{H}}^{U}$ built from a bypass decomposition of $M\setminus\big(\nu(K_{U})\cup V\big)$, although this need not be convex. Similarly, $\widehat{\mathcal{H}}^{V}$ is constructed from a bypass decomposition of $M\setminus\big(U\cup\nu(K_{V})\big)$. The relative bridges will be used in Proposition 3.10.
@@
[…]
@@
In order to construct a bridge splitting from a bypass decomposition, all the bypass attaching arcs must be isotoped off the faces of an $2$-handles already attached, but there is no canonical way to perform this isotopy. Although different isotopies preserve the weak contact isotopy type of each $U_{i}$, they may produce handlebodies which are inequivalent when the 2-handles are removed. We will show that an isotopy of the attaching 0-sphere for a $1$-handle on $\Sigma_{i}$ preserves the positive stabilisation class of the bridge $\widehat{\mathcal{H}}$.
Figure 7. Left: Before constructing the bridge of $\mathcal{H}$, bypass arcs must be isotoped off the faces of the 22-handles. Centre: Isotoping $c$ across the face of a $2$-handle on $\Sigma_{i}$ preserves the weak contact isotopy class of $U_{i}$. Right: When $h_{i}^{2}$ is removed, the bypass arcs $c^{R}$ and $c^{L}$ are not isotopic on $\widehat{\Sigma}$.
@@
Figure 9 shows that attaching the bypass along $c_{T}$ onto $\widehat{U}^{L}$ produces a handlebody weakly contact isotopic to $\widehat{U}^{R}$. Then by Theorems 2.7 and 3.4, the convex Heegaard decompositions $\widehat{\mathcal{H}}^{L}$ and $\widehat{\mathcal{H}}^{R}$ admit a common positve stabilisation. ∎
[…]
@@
Given a smooth Heegaard splitting $\mathcal{H}=(\Sigma,U,V)$ let $K_{U}$ and $K_{V}$ be Legendrian skeletons for $U$ and $V$ and let $X=M\setminus\big(\nu(K_{U})\cup\nu(K_{V})\big)$. Let $X_{1}\ast\dots\ast X_{f}$ and $X^{\prime}_{1}\ast\dots\ast X^{\prime}_{f^{\prime}}$ be distinct bypass decompositions of $X$, and let $\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(X_{1}\ast\dots\ast X_{f})$ and $\widehat{\mathcal{H}}^{\prime}=\widehat{\mathcal{H}}^{\prime}(X^{\prime}_{1}\ast\dots\ast X^{\prime}_{f})$ be the associated bridge Heegaard splittings formed by attaching only the $1$-handles of the respective bypasses to $\nu(K_{U})$. Then $\widehat{\mathcal{H}}$ and $\widehat{\mathcal{H}}^{\prime}$ are related by a sequence of positive (de)stabilisations.
@@
[…]
We claim that regardless of where the trivial bypass was inserted, $\widehat{\mathcal{H}}$ and $\widehat{\mathcal{H}}_{T}$ admit a common positive stabilisation. Recall that Proposition 3.8 allows us to isotope each attaching sphere along respective components of $\Gamma_{\Sigma_{i}}$ while preserving the positive stabilisation class of the bridge splitting. Using this flexibility, ensure that no handle attached after $h_{T}^{1}$ has an attaching sphere that intersects the disc supporting the trivial bypass. Not only can any attaching sphere be isotoped off this region, but we may also impose the stronger condition that subsequent attaching spheres are disjoint from the larger disc shown in Figure 10 that contains the bigon certifying $c_{T}$ as a trivial bypass arc. As a consequence, $c_{T}$ remains a trivial bypass arc on $\widehat{\Sigma}$; we will use this below. It follows that we may apply a sequence of Far Commutation moves to get a new bypass decomposition in which the bypass along $c_{T}$ is attached last.
@@
[…]
The relative bridge splitting $\widehat{\mathcal{H}}^{U}$ is tight by construction, so one may flip this Heegaard splitting upside down and apply Proposition 3.10 again to get the following consequence:
@@
The proof of Proposition 3.10 is essentially the same argument as the proof of [16, Theorem 6.1.]. Since it was not stated in full generality there, we indicate how it follows from the techniques of [16].
@@
Define the sequence of intermediate Heegaard splittings $\widehat{\mathcal{H}}_{j}=(\widehat{\Sigma}_{j},\widehat{U}_{j},\widehat{V}_{j})$ built by attaching only the first $j$ of the $1$-handles:
@@
[…]
Recall that the refinement is constructed by drilling tunnels through meridional discs for the handlebodies. We may choose these discs to be disjoint from the bypass $X_{1}=h^{1}_{1}\cup h^{2}_{1}$, so adding the $1$-handle $h_{1}^{1}$ to $U$ commutes with this drilling. Thus attaching $h^{1}_{1}$ to the refined $\widetilde{U}$ produces the same Heegaard splitting as taking the refinement of $\widehat{\mathcal{H}}_{1}$. As noted in Section 3.1, attaching $h_{1}^{1}$ in to $\widetilde{U}$ in the convex refinement is a positive stabilisation, which completes the base case.
Now proceed by induction on $j$. For the inductive step, note that if $\widehat{\mathcal{H}}_{j}$ is tight, then $\widehat{\mathcal{H}}_{j+1}$ is also tight because $\widehat{V}_{j+1}$ is a proper subset of the tight $\widehat{V}_{j}$, while $\widehat{U}_{j+1}$ is built from the tight $\widehat{U}_{j}$ by adding a contact $1$-handle. The argument above shows that refining and then adding a $1$-handle produces the same splitting as first adding the $1$-handle and then refining, which preserves the positive stabilisation class. This establishes the inductive step. ∎
@@ 3.3. Proof of the Giroux Correspondence
First, we repeat an argument from the proof of Theorem 6.1 in [16] that lets us restrict to Heegaard splittings that are smoothly isotopic:
@@
Choose Legendrian bouquets $K_{*}$ as the skeletons for each handlebody, where $\in\{U,U^{\prime},V,V^{\prime}\}$. By Theorem 2.1, after possibly some stabilistaions we may assume that the links $K_{U}\cup K_{V}$ and $K_{U^{\prime}}\cup K_{V^{\prime}}$ are Legendrian isotopic.
@@
Concatenating these produces a bypass decomposition of $X=M\setminus\big(\nu(K_{U})\cup\nu(K_{V})\big)$. As usual, let $\widehat{U}$ denote the handlebody formed by attaching all of the $1$-handles to $\nu(K_{U})$. Setting $)\widehat{\mathcal{H}}=(\partial\widehat{U},\widehat{U},M\setminus\text{int}(\widehat{U}$)), Corollary 3.11 implies that the refinement $\widetilde{\mathcal{H}}$ and the bridge $\widehat{\mathcal{H}}$ admit a common positive stabilisation. To complete the argument, we recall Remark 3.5: since $\mathcal{H}$ is convex, any refinement of $\widetilde{\mathcal{H}}$ is a positive stabilisation of $\mathcal{H}$ itself. Thus $\mathcal{H}$ and $\widehat{\mathcal{H}}$ admit a common positive stabilisation.